The demand for a certain product is 2000 units per year and the items are withdrawn at a constant rate. The ordering cost incurred each time an order is placed to replenish inventory is £50. The unit cost of purchasing the product is £470 per item, and the holding cost is £4.10 per item per year.

Apply a basic inventory model to determine the optimal size of each order and how often an order should be placed. You should follow the following steps:

(a) Formulate the mathematical problem.

(b) Determine the optimal size of each order.

(c) Determine how often an order should be placed. 


1
Expert's answer
2021-05-07T11:47:05-0400

Let q= numbers of unit ordered at a time

  

Total cost = commodity cost + ordering cost + holding cost


  "T=C.C.+O.C.+h.c."

     

  Commodity cost (CC)= units per year \times cost per unit 

            "= 2000\times 470=940000"

 

  Ordering cost "(OC)=\dfrac{50\times 2000}{Q}"


  Holding cost (HC) = cost of carrying"\times" average inventory


           "= 4.1\times \dfrac{Q+D}{2}"


 (a) "TC= 2000\times +\dfrac{50\times 2000}{Q}+\dfrac{4.1Q}{2}"


     "= 94000+\dfrac{100000}{Q}+2.05Q"


(b) For optimum size "\dfrac{d(TC)}{dq}=0"


           "\dfrac{d}{dq}(94000+\dfrac{10000}{q}+2.05q)=0"


          "\Rightarrow 0-\dfrac{100000}{q^2}+2.05=0"


           "q=\sqrt{\dfrac{100000}{2.05}}=220 \text{ units }"


(c) Length of production run/ or length of order "=\dfrac{1}{\text{ NO. of order}}=\dfrac{1}{\frac{U}{Q}}"


           "=\dfrac{Q}{U} year=\dfrac{12\times 20.86}{2000}=1.325 \text{ months }"


  Hence order shouls be placed after 1.325 months.